OpenStudy (anonymous):

Differential Equations: Find the solution for the given initial value problem. y''+y'-2y=2t y(0)=0 y'(0)=0

OpenStudy (anonymous):

I have the general solution as \[y=c _{1}e ^{-2t}+c _{2}e ^{t}\]

OpenStudy (anonymous):

Do you have the particular solution?

OpenStudy (anonymous):

no thats what i need ot get and im not sure how

OpenStudy (anonymous):

Use the method of undetermined coefficients

OpenStudy (anonymous):

It'll likely be related to the right side of the equal sign.

OpenStudy (anonymous):

i know that im just not sure what to define as \[Y(t)\]

OpenStudy (anonymous):

See http://tutorial.math.lamar.edu/Classes/DE/UndeterminedCoefficients.aspx for some good notes and examples.

OpenStudy (anonymous):

Try Y=At+B

OpenStudy (anonymous):

@CliffSedge yes

OpenStudy (anonymous):

The basic method goes: \[y''+y'-2y = 2t\] we assume that y = at+b (in this case); \[0+a-2(at+b) = 2t\]

OpenStudy (anonymous):

Find A and B by taking first and second derivatives of At+B and plug into original equation.

OpenStudy (anonymous):

well i got Y(t)= 2At Y'(t)=2A Y''(t)=2A?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

you have: \[-2at = 2t\] and \[-2b + a = 0\]

OpenStudy (anonymous):

you can cancel out the t's in the first equation: \[-2a = 2\] then solve for a and b

OpenStudy (anonymous):

ok i understand but how did you get -2at=2t and -2b+a=0

OpenStudy (anonymous):

Y = At+B Y' = A Y'' = 0

OpenStudy (anonymous):

so lets go back to the equation we had previously: \[a -2(at+b) = 2t\] so: \[(-2at)+(a-2b) =(2t)+(0)\]

OpenStudy (anonymous):

oh i see now ok i got it from here thank you very much

OpenStudy (anonymous):

the coefficients of t^1 must equal, and the coefficients of t^0 must equal

OpenStudy (anonymous):

you are welcome